The specific heats of some elements at 25°C are as follows: aluminum = 0.215 cal/g⋅°C; carbon (graphite) = 0.170 cal/g⋅°C; iron = 0.107 cal/g⋅°C; mercury = 0.0331 cal/g⋅°C.

a) Which element would require the smallest amount of heat to raise the temperature of 100. g of the element by 10.°C?

  • iron
  • mercury
  • aluminum
  • carbon (graphite)

b) If the same amount of heat needed to raise the temperature of 1 g of aluminum by 59°C were applied to 1 g of mercury, by how many degrees would its temperature be raised?

ΔTHg = ..........°C

c) If a certain amount of heat is used to raise the temperature of 1.6 g of iron by 10.°C, the temperature of 1 g of which element would also be raised by 10.°C, using the same amount of heat?

  • aluminum
  • carbon (graphite)
  • mercury
1
Expert's answer
2021-01-25T04:20:04-0500

1. Q=mc∆T.

∆T= Q/mc. Therefore it is aluminium.

2. Q= 1 x 0.215 x 59 = 12,685.

∆T = 12,685/1 x 0.0331 = 383,232628.

3. Q= 1.6 x 0.107 x 10 = 1,712.

c= Q/m∆T = 1,712/ 1 x 10 = 0,1712, it is carbon.